PERTURBATIONS OF THE FINE AND HYPERFINE STRUCTURE IN THE PURE ROTATIONAL SPEC-TRUM OF VCl (X^{5}Δ_{r})

\underline{D. T. HALFEN} and L. M. ZIURYS, Department of Chemistry, Department of Astronomy, and Steward Observatory, University of Arizona, Tucson, AZ, 85721.

The pure rotational spectrum of the VCl (X^{5}Δ_{r}) radical has been measured using millimeter/sub-mm-wave direct absorption methods. This work is the first pure rotational study of this molecule. Vanadium chloride was produced in a plasma created by an AC discharge of gas-phase VCl_{4} and argon. Ten rotational transitions each of the ^{35}Cl and ^{37}Cl isotopomers have been measured for all five spin-orbit components, confirming the ^{5}Δ_{r} ground state. Lambda doubling is observed in four of the five Ω ladders, including the Ω = 0 ladder, which had not been observed previously. Each fine structure component was found to exhibit hyperfine splittings due to the ^{51}V nuclear spin (I = 7/2). The fine structure pattern of the spin-orbit ladders suggests that the ground state is perturbed with the Ω = 1 ladder significantly shifted with respect to the overall pattern. In addition, the hyperfine structure of the Ω = 1 and 2 ladders becomes highly irregular for certain rotational transitions. Rotational, spin-orbit, spin-spin, lambda-doubling, and hyperfine parameters have been determined from the spectra, and the results will be discussed.